40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 14 12:13:52 2009, Jenne, Update, Treasure, We are *ROCKSTARS* ! IFO is back up Jenne14Dec09_IFOlocked.png
    Reply  Mon Dec 14 13:08:33 2009, Koji, Update, Treasure, LOCKSTARS 
       Reply  Mon Dec 14 14:16:14 2009, Alberto, Update, Treasure, LOCKSTARS 
    Reply  Mon Dec 14 13:17:33 2009, rob, Update, Treasure, We are *ROCKSTARS* ! IFO is back up two-thumbs-up.jpeg
Message ID: 2410     Entry time: Mon Dec 14 12:13:52 2009     Reply to this: 2411   2412
Author: Jenne 
Type: Update 
Category: Treasure 
Subject: We are *ROCKSTARS* ! IFO is back up 

[Jenne, Kiwamu, Koji]

We got the IFO back up and running!  After all of our aligning, we even managed to get both arms locked simultaneously.  Basically, we are awesome. 

 This morning, we did the following:

*  Turned on the PZT High voltages for both the steering mirrors and the OMC.  (For the steering mirrors, turn on the power, then hit "close loop" on each.  For the OMC, hit Output ON/OFF).

*  Looked at the PZT strain gauges, to confirm that the PZTs came back to where they had been.  (Look at the snapshot of C1ASC_PZT_Al)

*  Locked all components of the PSL (This had already been done.)

*  Removed beam dump which was blocking the PSL, and opened the PSL mechanical shutter.  Light into the IFO!

*  Locked the Mode Cleaner.  The auto-locker handled this with no problem.

*  Confirm that light is going through the Faraday.  (Look at the TV sitting on top of MC13 tank...it shows the Faraday, and we're hitting the input of the Faraday pretty much dead-on).

*  Look at IP_ANG and IP_POS.  Adjust the steering mirrors slightly to zero the X&Y readings on IP_ANG.  This did not change the PZTs by very much, so that's good.

*  Align all of the Core Optics to their OpLev positions.

*  On the IFO_Align screen, save these positions.

*  Run the IFO_Configure scripts, in the usual order.  (Xarm, Yarm, PRM, DRM).  Save the appropriate optics' positions after running the alignment scripts.  We ended up running each alignment script twice, because there was some residual misalignment after the first iteration, which we could see in the signal as viewed on DataViewer (Either TRX, TRY, or SPOB, for those respective DoFs).

*  Restore Full IFO.

*  Watch the beauty of both arms and the central cavity snapping together all by themselves!  In the attached screenshot, notice that TRX and TRY are both ~0.5, and SPOB and AS166Q are high.  Yay!

Conclusions: 

*  The wiping may have helped.  While aligning X and Y separately, TRX got as high as ~1.08, and TRY got as high as 0.98  This seems to be a little bit higher than it was previously.

*  Since everything locked up in pretty short order, and the free swinging spectra (as measured by Kiwamu in elog 2405) looks good, we didn't break anything while we were in the chambers last week.  Excellent.

*  We are now ready for a finesse measurement to tell us more quantitatively how we did with the wiping last week.

 

Attachment 1: Jenne14Dec09_IFOlocked.png  417 kB  | Hide | Hide all
Jenne14Dec09_IFOlocked.png
ELOG V3.1.3-